urmum:

Solve for w. 5w + 9z = 2z + 3w

Astrid1:

So this is our equation: 5w + 9z = 2z + 3w Subtract 9z from each side. \[5w+9z-9z=2z+3w-9z\] Simplify: \[5w=-7z+3w\] Subtract 3w from each side. \[5w-3w=-7z+3w-3w\] Simplify: \[2w=-7z\] Divide both sides by two. \[\frac{2w}{2}=\frac{-7z}{2}\] Simplify. \[w=-\frac{7z}{2}\]

Astrid1:

@rryyyyyyyyyy wrote:
but it the answer bud
I know, but it's against the rules to give away a direct answer without an explanation.